Peat is gunning the engine on his sports car. He finds that he can increase the engine RPM, or rotations per minute, from 2000 to 5000 in about a third of a second. What is the average angular acceleration of the engine shaft during this time interval? Select one: a. 56,000 rad / s2 b. 940 rad / s2 c. 310 rad / s2 d. 150 rad / s2
Answer #1

Initial rpm, w1 = 2000 rpm = 2000*2pi/60 = 209.33 rad/s
Final rpm, w2 = 5000 rpm = 5000*2pi/60 = 523.33 rad/s
Avg angular acc = (w2 - w1)/t = (523.33-209.33)/0.33 = 940 rad/s/s